21 Milano Co. manufactures and sells three products: I, product 2. and product 3. Their unit selling prices are product 1, $40: product 2. S30, and product 3, S20. The per unit variable costs to manufacture and sell these products are product 1 S30; product 2, S15; and product 3.SS Their sales mix is reflected in a ratio of 6:4:2 Annual fixed costs shared by all three products are $270.000. One type of raw material has been used to manufacture products 1 and 2. The company has developed a new material of equal quality for less cost. The new material would reduce variable costs per unit as follows: product 1 by $10 and product 2 by S5. However, the new material requires new equipment, which will increase annual fixed costs by S50,000. Required 1. If the company continues to use the old material, determine its break-even point in both sales units and sales dollars of each individual product. Check (f) Old plan break-even, 1,875 composite units 2 If the company uses the new material, determine its new each individual product. (Round to the next whole unit.) break-even point in both sales units and sales dollars of 2
alip Exercise 21-17 Target income and margin of safety (in dollars) c20 Refer to the information in Exercise 21-16 as a target pretax income of $162.000 for 2013. What amount of sales (in dollars) is needed to produce this target income? 2. If Hudson achieves its target pretax income for 2013, what is its margin of safety (in percent)? (Round to one decimal place.) HUDSON CO. Contribution Margin Income Statement For Year Ended December 31, 2017 Sales (9.600 units at S225 each) Variable costs (9,600 units at $180 each) 1.25000.2 Contribution margin $2.160,000 432,000 324.000 S 108,000 loPretax income (e Exercise 21-25 Degree of operating leverage A2e Refer to the information in Exercise 21-16. 1. Compute the companys degree of operating leverage for 2017 If sales decrease by 5% in 2013, what will be the companys pretax income? 3, Assume sales for 20 is decrease by 5%. Prepare a contribution margin income statement for 2013.
Answer #1

Answer

  • All working forms part of the answer
  • Exercise 21 – 78

--Requirement 1

Working

Product 1

Product 2

Product 3

A

Sale price per unit

$                40.00

$                 30.00

$          20.00

B

Variable Cost per unit

$                30.00

$                 15.00

$            8.00

C = A - B

Contribution margin per unit

$                10.00

$                 15.00

$          12.00

D

Sales Mix

0.5

0.333333333

0.16666667

E = C x D

Weighted Contribution margin per unit

$                   5.00

$                    5.00

$            2.00

A

Annual Fixed Cost

$      270,000.00

B = $ (5+5+2)

Total Weighted average contribution margin per unit

$                12.00

C = A/B

Break Even point in Sales unit

22500

Answer

Working

Product 1

Product 2

Product 3

A

Sales Mix

0.5

0.333333333

0.16666667

B = 22500 units x A

Individual Break Even in units - ANSWER

11250

7500

3750

C

Sale price per unit

$                40.00

$                 30.00

$          20.00

D = B x C

Individual Break Even in Sales Dollars – ANSWER

$      450,000.00

$       225,000.00

$ 75,000.00

---Requirement 2

Working

Product 1

Product 2

Product 3

A

Sale price per unit

$            40.00

$                30.00

$               20.00

B

Variable Cost per unit

$            20.00

$                10.00

$                 8.00

C = A - B

Contribution margin per unit

$            20.00

$                20.00

$               12.00

D

Sales Mix

0.5

0.333333333

0.166666667

E = C x D

Weighted Contribution margin per unit

$            10.00

$                  6.67

$                 2.00

A

Annual Fixed Cost

$ 270,000.00

B = $ (5+5+2)

Total Weighted average contribution margin per unit

$            18.67

C = A/B

Break Even point in Sales unit

14464

Answer

Working

Product 1

Product 2

Product 3

A

Sales Mix

0.5

0.333333333

0.166666667

B = 22500 units x A

Individual Break Even in units - ANSWER

7232

4821

2411

C

Sale price per unit

$            40.00

$                30.00

$               20.00

D = B x C

Individual Break Even in Sales Dollars – ANSWER

$ 289,280.00

$     144,630.00

$      48,220.00

  • Exercise 21.17

--Working

A

Contribution margin

$            432,000

B

Sales

$        2,160,000

C = (A/B) x 100

CM ratio

20.0%

--Requirement 1

A

Target pretax Income

$            162,000

B

Fixed Cost

$            324,000

C= A+B

Total Contribution margin required

$            486,000

D

CM ratio

20.0%

E = C/D

Sales Dollars required

$        2,430,000 = Answer

---Requirement 2

A

Actual Sale for 2018

$        2,430,000

B

Fixed Cost

$            324,000

C

CM ratio

20.0%

D = B/C

Break Even Sales Dollars

$        1,620,000

E = A - D

Margin of Safety Sales dollars

$            810,000

F = (E/A) x 100

Margin of Safety %

33.3% = Answer

  • Exercise 21-25

--Requirement 1

A

Contribution margin

$            432,000

B

Pretax Income

$            108,000

C = A/B

Degree of Operating Leverage

4 = Answer

--Requirement 2

A

Decrease in Sales %

5%

B

Degree of Operating Leverage

4

C = A x B

Decreases in Pretax Income %

20.0%

D

2017's Pretax Income

$            108,000

E = D x C

Decrease in 2018

$              21,600

F = D - E

Actual Pretax Income for 2018

$              86,400 = Answer

--Requirement 3

Income Statement for 2018

Sales [9120 units at $225]

$        2,052,000

Variable Costs [9120 units x $ 180]

$        1,641,600

Contribution margin

$            410,400

Fixed Cost

$            324,000

Pretax Income [2018]

$              86,400